OpenStudy (crashonce):

If the sum to 'n' terms of an AP is n^2 + 2n, find the 6th term of the sequence.

OpenStudy (crashonce):

@ganeshie8 sorry

ganeshie8 (ganeshie8):

Consider an arithmetic progression : \[\rm \large a_1, a_2, a_3, a_4, a_5, a_6, \ldots\]

ganeshie8 (ganeshie8):

if you add the first 6 terms, then subtract the sum of first 5 terms, what do you get ?

ganeshie8 (ganeshie8):

Say : sum of first 6 terms is \(\rm S_6\) sum of first 5 terms is : \(\rm S_5\)

ganeshie8 (ganeshie8):

can we say \[\rm S_6 - S_5 = a_6\] ?

OpenStudy (crashonce):

uh yea

ganeshie8 (ganeshie8):

find the value of \(\rm S_6\), \(\rm S_5\) and plugin above

ganeshie8 (ganeshie8):

use the given formula to find the sums

ganeshie8 (ganeshie8):

\[\rm\begin{align} S_n &= n^2+2n\\~\\S_6&= ?\\~\\S_5&= ?\end{align}\]

OpenStudy (crashonce):

s6 = 36+12 = 48 s5 = 25+10 = 35

OpenStudy (crashonce):

ooooo its 13 right thanks

ganeshie8 (ganeshie8):

Excellent !